By method variation of parameter solve the ODE(D^2+4)y=2 cosec2x
1
Expert's answer
2021-07-29T16:35:36-0400

"\left( {{D^2} + 4} \right)y = y'' + 4y = 2{\rm{cosec}}2x"

Characteristic equation:

"{k^2} + 4 = 0 \Rightarrow {k^2} = - 4 \Rightarrow k = \pm 2i"

Then the general so lution of the homogeneous equation is

"{y_0} = {C_1}\cos 2x + {C_2}\sin 2x"

Let

"{C_1} = f(x),\,\,{C_2} = g(x)"

Then

"y = f\cos 2x + g\sin 2x \Rightarrow y' = f'\cos 2x - 2f\sin 2x + g'\sin 2x + 2g\cos 2x"

Let

"f'\cos 2x + g'\sin 2x = 0\,\,\,\,(*)"

Then

"y' = - 2f\sin 2x + 2g\cos 2x \Rightarrow y'' = - 2f'\sin 2x - 4f\cos 2x + 2g'\cos 2x - 4g\sin 2x"

Substitute the found values ​​into the original equation:

"- 2f'\sin 2x - 4f\cos 2x + 2g'\cos 2x - 4g\sin 2x + 4f\cos 2x + 4g\sin 2x = 2{\rm{cosec}}2x"

"- 2f'\sin 2x + 2g'\cos 2x = 2{\rm{cosec}}2x"

Together with equation (*) we have the system of equations:

"\left\{ \begin{array}{l}
 - 2f'\sin 2x + 2g'\cos 2x = 2{\rm{cosec}}2x\\
f'\cos 2x + g'\sin 2x = 0
\end{array} \right."

From the second equation we obtain

"g' = - f' \cdot \frac{{\cos 2x}}{{\sin 2x}}"

Substitute the found value into the first equation:

"- 2f'\sin 2x + 2\left( { - f' \cdot \frac{{\cos 2x}}{{\sin 2x}}} \right)\cos 2x = 2{\rm{cosec}}2x"

"- 2f'\sin 2x - 2f' \cdot \frac{{{{\cos }^2}2x}}{{\sin 2x}} = \frac{2}{{\sin 2x}}"

"- 2f' \cdot \frac{{{{\sin }^2}2x + {{\cos }^2}2x}}{{\sin 2x}} = \frac{2}{{\sin 2x}}"

"\frac{{ - 2f'}}{{\sin 2x}} = \frac{2}{{\sin 2x}}"

"f' = - 1 \Rightarrow f = - x + {C_3}"

"g' = - f' \cdot \frac{{\cos 2x}}{{\sin 2x}} = \frac{{\cos 2x}}{{\sin 2x}} \Rightarrow g = \int {\frac{{\cos 2x}}{{\sin 2x}}dx} = \frac{1}{2}\int {\frac{{d\sin 2x}}{{\sin 2x}} = \frac{1}{2}} \ln |\sin 2x| + {C_4}"

Then

"y = f\cos 2x + g\sin 2x = \left( { - x + {C_3}} \right)\cos 2x + \left( {\frac{1}{2}\ln |\sin 2x| + {C_4}} \right)\sin 2x"

Answer: "y =\left( { - x + {C_3}} \right)\cos 2x + \left( {\frac{1}{2}\ln |\sin 2x| + {C_4}} \right)\sin 2x"